find the greatest six digit  number that exactly divisible by 27 45 60 72 and 96

Grade:12th pass

1 Answers

Gulshan Goyal
6 years ago
lcm of given numbers is 4320
greatest 6 digit no. can be calculated as  [1000000/4320]=231,
Answer=231*4320=997920